OpenStudy (anonymous):

Write an equation and a solution for the problem. A health food store sells oatmeal for $3.60 per pound and bran flakes for $4.80 per pound. How many pounds of each should be used to get a mixture of 30 pounds that sells for $4.00 a pound?

OpenStudy (anonymous):

killing them with these mixture problem, huh?

Parth (parthkohli):

I hate these >.<

OpenStudy (anonymous):

lets cut to the chase \[4.80x+3.60(30-x)=30\times 4\] \[48x+36(30-x)=1200\] http://www.wolframalpha.com/input/?i=48x%2B36%2830-x%29%3D1200 10 pounds of the $4.80 mixture and so 20 of the $3.60